OpenStudy (anonymous):

integral 1 to 2 ((x^2-x)/(x^3))

geerky42 (geerky42):

Let me rewrite it so it can be read easily: \[\large \int\limits_{1}^{2} \dfrac{x^2-x}{x^3}dx\]

geerky42 (geerky42):

ah ok, you can split the function like this: \[\large \int_1^2 \dfrac{x^2}{x^3} - \dfrac{x}{x^3} dx = \int_1^2 \dfrac{1}{x} - \dfrac{1}{x^2} dx \\ \large = \int_1^2 \dfrac{1}{x}dx - \int_1^2 \dfrac{1}{x^2} dx\]

geerky42 (geerky42):

Does that help you?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

does the 1/x become ln?

geerky42 (geerky42):

yes

OpenStudy (anonymous):

can you help me with another integral question plz?

geerky42 (geerky42):

Sure

OpenStudy (anonymous):